2x -4y=8
x -2y=4 trying to find out how to get the value of x and y

Answer by stanbon(57377)

2x -4y=8
x -2y=4
-----------------
Multiply thru the 2nd equation by 2 to get:
2x-4y = 8
2x-4y = 8
----
Subtract to get: 0 = 0
-----
The two equations are really just one equation.
The solution is every point on the line
y = (1/2)x-2
or all points of the form (x, (1/2)x-2)
=============================================
Cheers,
stan H.
=================